Marcus has a pension plan with an annuity that is guaranteed to earn 13% interest compounded semi-annually. He plans to work for 10 years before retiring and would then like to be able to draw an income of $110,000.00 per annum for 18 years. How much should he deposit annually into his retirement fund to accomplish this?


1
Expert's answer
2022-02-28T09:39:14-0500

PV=110 000

n=36 ( 18*2)

r=0.065

the current (present) value of an ordinary annuity

"PV=A(\frac{1-\frac{1}{(1+r)^n}}{r})"


"110 000=A(\frac{1-\frac{1}{(1+0.065)^{36}}}{0.065})"


A (0.5)=7976.47


A(year)=15952.93